OpenStudy (anonymous):

sqrt(4x + 16) = x + 1

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

take square on both sides

OpenStudy (anonymous):

yes i got that. This is what I have so far: square both sides 4x+16=x^2+0 4x+16=x^2

OpenStudy (anonymous):

You made a mistake

OpenStudy (anonymous):

which part?

OpenStudy (anonymous):

it would be x^2+2x+1

OpenStudy (anonymous):

for the right side

OpenStudy (anonymous):

oh oops that is rigt :P i just typed it wrong lol thank you

OpenStudy (anonymous):

So, what would you do next?

OpenStudy (anonymous):

I dont know. I'm stuck after that. That's why i'm asking for help :P

OpenStudy (anonymous):

shift x on left side and constant values on right

OpenStudy (anonymous):

You should rearrange to get x^2+2x+1=4x+16, Then x^2-2x-15=0

OpenStudy (anonymous):

Then factor that last equation.

OpenStudy (anonymous):

(x+3)(x-5)=0

OpenStudy (anonymous):

4x+16=x^2+2x+1 4x-2x-x^2=1-16 -X^2+2x=-15 x^2-2x=15 x^2-2x+15 now made factors

OpenStudy (anonymous):

@03225186213 You made a little mistake towards the end. it would be x^2-2x-15=0. Otherwise that's correct.

OpenStudy (anonymous):

oops thanks

OpenStudy (anonymous):

(x+3)(x-5)=0 x=-3 x=5